# The Devil is in the EOS: Sequence Training for Detailed Image Captioning

### v1

@@ 1 Introduction
As vision-language models (VLMs) advance, new frontiers open up in multimodal research, some examples being the novel tasks of interactive visual dialogue (Liu et al., 2023) and culturally-aware visual question answering (Romero et al., 2024). The fundamental task of image captioning (Vinyals et al., 2015) has also seen an upgrade: in place of the classic benchmark MS COCO (Lin et al., 2014), a number of new evaluation datasets have been proposed, consisting of images with detailed (a.k.a fine-grained or dense) captions (Onoe et al., 2024; Cho et al., 2022; Urbanek et al., 2024). Detailed image captioning has extrinsic applications, e.g., assistive applications, as well as a key intrinsic application, as a form of image textualization for solving knowledge- and reasoning-intensive vision-language tasks using large language models (Pi et al., 2024).
@@
We view the generation of detailed captions as a form of recovering existing model capabilities (rather than instilling new ones) and propose to achieve that through end-of-sequence (EOS) token debiasing. The EOS token exists in every single caption in the training data; combined with teacher forcing, this causes the model to have a bias towards predicting it. Elevating this bias should lead the model to generate longer text (Newman et al., 2020; Kulikov et al., 2021), and the conditioning on the image context during vision-language pretraining should ensure that this longer text remains grounded in the input image. Through our experiments, we show that simply pushing down the probability of the EOS token through sequence training of models trained on generic captions leads to longer, more detailed and largely accurate outputs.
@@ 3.1 Sequence Training for Image Captioning
Rennie et al. (2017) adapt sequence training to image captioning, replacing the traditional next-token prediction objective with direct optimization of an evaluation metric applied to complete sequences (captions) generated by the model. Given an input image $I$, a sequence of tokens is generated, $S=[t_{1},t_{2},....,t_{n}]$ where $t_{n}$ is the EOS token. Next, a reward, $R(S)$, is calculated against the reference captions for $I$ using the CIDEr metric (Vedantam et al., 2015). Finally, the gradient is computed using REINFORCE (Zhang et al., 2020):
@@ 4 Experiments
We test the proposed approach on three VLMs, built on strong language backbones and finetuned for the task of image captioning: BLIP-2 with an OPT 2.7B decoder, [^2] Salesforce/BLIP-2-opt-2.7b-coco BLIP-2 with a FlanT5-XL decoder, [^3] Salesforce/BLIP-2-flan-t5-xl-coco and PaliGemma 3B444google/paligemma-3b-ft-cococap-224. We choose the checkpoints finetuned on the MS COCO dataset (Lin et al., 2015) and adopt the same dataset in our experiments, as this ensures a good distribution alignment between the initial finetuning and our sequence training, and allows us to quantify exactly how much extra detail we can induce through EOS debiasing for the same images that the base model was trained on. The full recipe we propose essentially amounts to two-stage training on the same data: finetuning with the standard next-token prediction objective, and sequence training with EOS debiasing to generalize beyond the length learned in the first stage.
@@ 4.1 Implementation details
We finetune only the cross-modal bridge of the VLM (Q-former and linear projection in BLIP-2, linear layer in PaliGemma). The rest of the model parameters are frozen. The learning rate is set to $1e$-777, to allow the VLM to gradually find a solution to the longer generation task. We find that higher values lead to degenerate model behavior, while lower values excessively delay model convergence. Learning rate ablation can be found in A.3.2. Models are trained with a batch size of 8 and 3 gradient accumulation steps on an NVIDIA A100 80GB card. We use contrastive decoding (Su & Collier, 2022) during training, to encourage exploration through more diverse outputs. We train until the generated captions reach a sequence length of 60, the maximum that fits on this GPU card. All generation is done using beam search decoding with 555 beams, a repetition penalty of 1.51.51.5 and a no-repeat-ngram constraint of 333.
@@
Table 1: Summary of the evaluation datasets in terms of number of images we use, number of captions available per image, average length (in words) per caption, and additional details.
@@ Acknowledgments
We thank Jameel Hassan for his valuable discussions and insights during the early stages of this work.
@@ A.1 Extra Details
[Uncaptioned image] References A man in orange jersey throwing a baseball. ||| A young baseball player is on the field with a mitt. ||| A baseball pitcher steps forward as he winds up a pitch. ||| A pitcher is winding up for a pitch. ||| A pitcher in mid pitch at a baseball game.
@@ A.3.1 Trivial baseline
[…]
@@ A.4 GPT-4o Prompt
Table 9 shows the prompt used to induce Coherence scores from GPT-4o. We use a temperature of 1.01.01.0 when calling the API.

### v2

@@ 1 Introduction
As vision-language models (VLMs) advance, new frontiers open up in multimodal research, some examples being the novel tasks of interactive visual dialogue (Liu et al., 2023) and culturally-aware visual question answering (Romero et al., 2024). The fundamental task of image captioning (Vinyals et al., 2015) has also seen an upgrade: in place of the classic benchmark MS COCO (Lin et al., 2014), a number of new evaluation datasets have been proposed, consisting of images with detailed (a.k.a fine-grained or dense) captions (Onoe et al., 2024; Cho et al., 2022; Urbanek et al., 2024). Detailed image captioning has extrinsic applications, e.g., assistive tools, as well as a key intrinsic application, as a form of image textualization for solving knowledge- and reasoning-intensive vision-language tasks using large language models (Pi et al., 2024).
@@
We view the generation of detailed captions as a form of recovering existing model capabilities (rather than instilling new ones) and propose to achieve that through end-of-sequence (EOS) token debiasing. The EOS token exists in every single caption in the training data; combined with teacher forcing, this causes the model to have a bias towards predicting it. Alleviating this bias should lead the model to generate longer text (Newman et al., 2020; Kulikov et al., 2021), and the conditioning on the image context during vision-language pretraining should ensure that this longer text remains grounded in the input image. Through our experiments, we show that simply pushing down the probability of the EOS token through sequence training of models trained on generic captions leads to longer, more detailed and largely accurate outputs.
@@ 3.1 Sequence Training for Image Captioning
Rennie et al. (2017) adapt sequence training to image captioning, replacing the traditional next-token prediction objective with direct optimization of an evaluation metric applied to complete sequences (captions) generated by the model. Given an input image $I$, a sequence of tokens is generated, $S=[t_{1},t_{2},....,t_{n}]$ where $t_{n}$ is the EOS token. Next, a reward, $R(S)$, is calculated against the reference captions for $I$ using the CIDEr metric (Vedantam et al., 2015). Finally, the gradient is computed using REINFORCE (Williams & Zipser, 1989):
@@ 4 Experiments
We test the proposed approach on three VLMs, built on strong language backbones and finetuned for the task of image captioning: BLIP-2 with an OPT 2.7B decoder, [^2] Salesforce/BLIP-2-opt-2.7b-coco BLIP-2 with a FlanT5-XL decoder, [^3] Salesforce/BLIP-2-flan-t5-xl-coco and PaliGemma 3B444google/paligemma-3b-ft-cococap-224. We choose the checkpoints finetuned on the MS COCO dataset (Lin et al., 2014) and adopt the same dataset in our experiments, as this ensures a good distribution alignment between the initial finetuning and our sequence training, and allows us to quantify exactly how much extra detail we can induce through EOS debiasing for the same images that the base model was trained on. The full recipe we propose essentially amounts to two-stage training on the same data: finetuning with the standard next-token prediction objective, and sequence training with EOS debiasing to generalize beyond the length learned in the first stage.
@@ 4.1 Implementation details
We finetune only the cross-modal bridge of the VLM (Q-former and linear projection in BLIP-2, linear layer in PaliGemma). The rest of the model parameters are frozen. The learning rate is set to 1e1e-77, to allow the VLM to gradually find a solution to the longer generation task. We find that higher values lead to degenerate model behavior, while lower values excessively delay model convergence. Learning rate ablation can be found in A.3.2. Models are trained with a batch size of 8 and 3 gradient accumulation steps on an NVIDIA A100 80GB card. We use contrastive decoding (Su & Collier, 2022) during training, to encourage exploration through more diverse outputs. We train until the generated captions reach a sequence length of 60, the maximum that fits on this GPU card. All generation is done using beam search decoding with $5$ beams, a repetition penalty of $1.5$ and a no-repeat-ngram constraint of $3$.
@@
Table 1: Summary of the evaluation datasets in terms of number of images we use, number of captions available per image, average length (in words) per caption, and other details.
@@ Acknowledgments
We thank Jameel Hassan for his contribution in the early stages of this work.
@@ A.1 Extra Details
[Uncaptioned image] References A man in orange jersey throwing a baseball. $|$ A young baseball player is on the field with a mitt. $|$ A baseball pitcher steps forward as he winds up a pitch. $|$ A pitcher is winding up for a pitch. $|$ A pitcher in mid pitch at a baseball game.
@@ A.3.1 Trivial baseline
For completeness, Table 7 compares the trivial solution under different decoding strategies: beam search, contrastive decoding (Su & Collier, 2022), and nucleus sampling (Holtzman et al., 2019), against our approach. The beam size for beam search is set to $5$, the contrastive decoding parameters are $\alpha=0.7$ and $k=5$, and the top-$p$ for nucleus sampling is $0.9$.
@@ A.4 GPT-4o Prompt
Table 9 shows the prompt used to induce Coherence scores from GPT-4o. We use a temperature of $1.0$ when calling the API.